OpenStudy (lukecrayonz):

Find the derivative of y=x^5lnx-1/4x^4

OpenStudy (lukecrayonz):

http://gyazo.com/a755e8a51204ef28a3006a202c67634e

OpenStudy (lukecrayonz):

@jim_thompson5910 :D?

zepdrix (zepdrix):

\[\Large y\quad=\quad x^5 \ln x-\frac{1}{4}x^4\]\[\Large y'\quad=\quad \color{royalblue}{(x^5)'}\ln x+ x^5\color{royalblue}{(\ln x)'}-\frac{1}{4}\color{royalblue}{(x^4)'}\]For the first term, we'll need to apply the product rule, yes? The blue parts represent derivatives we need to take.

OpenStudy (lukecrayonz):

5x^4lnx+x^5*1/x-1/4(4x^3)

OpenStudy (lukecrayonz):

@zepdrix

zepdrix (zepdrix):

Ya looks good so far! Now it just needs a little simplification.

OpenStudy (lukecrayonz):

x^4+5x^5ln(x)-x^3?

zepdrix (zepdrix):

\[\Large y'\quad=\quad x^4+5x^{\color{red}{5}}\ln(x)-x^3\]Hmm looks ok besides this one small part here.

OpenStudy (lukecrayonz):

5x^4lnx+x^4-x^3!

zepdrix (zepdrix):

yay good job!